Required information [The following information applies to the questions displayed below.] Actuary and trustee reports indicate the following changes in the PBO and plan assets of Lakeside Cable during 2018: Prior service cost at Jan. 1, 2018, from plan amendment at the beginning of 2016 (amortization: $6 million per year) $ 46 million Net loss–pensions at Jan.1, 2018 (previous losses exceeded previous gains) $ 64 million Average remaining service life of the active employee group 10 years Actuary’s discount rate 10 % ($ in millions) PBO Plan Assets Beginning of 2018 $ 440 Beginning of 2018 $ 300 Service cost 62 Return on plan assets, Interest cost, 10% 44 8.0% (10% expected) 24 Loss (gain) on PBO (4 ) Cash contributions 78 Less: Retiree benefits (52 ) Less: Retiree benefits (52 ) End of 2018 $ 490 End of 2018 $ 350 3. Prepare a pension spreadsheet to assist you in determining end of 2018 balances in the PBO, plan assets, prior service cost—AOCI, the net loss—AOCI, and the pension liability. (Enter your answers in millions (i.e., 10,000,000 should be entered as 10). Enter credit amounts with a minus sign and debit amounts with a positive sign.)
Answer #1

Part 3

PBO

Plan Assets

Prior Service Cost–AOCI

Net Loss–AOCI

Pension expense

Cash

Net pension (liability) / asset

Bal. Jan. 1, 2018

440

300

46

64

(140)

Service cost

(62)

62

(62)

Interest cost, 10%

(44)

44

(44)

Expected return on assets

30

(30)

30

Loss on assets

(6)

6

(6)

Amortization of:

Prior service cost–AOCI

(6)

6

Net loss–AOCI (64-(440*10%)/10

(2)

2

Gain on PBO

4

(4)

Cash contributions

78

(78)

78

Retiree benefits

52

(52)

Bal., Dec. 31, 2018

(490)

350

40

64

84

(144)

Expected return on assets = 24*10%/8% = 30

Loss on assets = 30-24 = 6